OpenStudy (anonymous):

Billy is analyzing a circle, y2 + x2 = 81, and a linear function g(x). Will they intersect? y2 + x2 = 81 g(x) graph of the function y squared plus x squared equals 81 x g(x) −1 −8.5 0 −9 1 −9.5 Yes, at positive x coordinates Yes, at negative y coordinates Yes, at negative and positive x coordinates No, they will not intersect @Directrix @Callisto

OpenStudy (anonymous):

@iPwnBunnies

Directrix (directrix):

I guess we have to get the equation of the line and then see if it intersects the circle.

OpenStudy (anonymous):

ok

OpenStudy (triciaal):

yes at

Directrix (directrix):

You can use the online calculator to generate the equation of the line through two given points. The calculator is at this link: http://www.mathportal.org/calculators/analytic-geometry/two-point-form-calculator.php

OpenStudy (anonymous):

@Callisto

Directrix (directrix):

If you want to use it, here is how the image there looks when you key in the data.

Directrix (directrix):

@Callisto My explanation does not resonate with Asker. Asker did not participate so I'll leave the discussion with you. :)

OpenStudy (anonymous):

i just dont get any of this stuff i'm sorry

Directrix (directrix):

No need to apologize for not getting this stuff. @iamabarbiegirl

OpenStudy (anonymous):

I just tryr to get help and no matter who i tag no one answers my questions and it just really stresses me out i really am sorry

OpenStudy (callisto):

If you don't have a graphing calculation, here is what you can do: Step 1. Find the equation of g(x). As given in the question, g(x) is a linear function. Since you are given at least two points, you can use the two point form: \(\frac{y-y_1}{x-x_1} = \frac{y_2-y_1}{x_2-x_1}\) to find equation. Maybe you can pick two points from the given three points first?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

I don't get thisss!!! D':

OpenStudy (callisto):

What don't you understand?

OpenStudy (anonymous):

how to get the answerr

OpenStudy (anonymous):

and everything else about this question

OpenStudy (callisto):

I've just told you the very first step. Which part of it do you not understand?

OpenStudy (anonymous):

i dont have time for a slow answer.

OpenStudy (anonymous):

i want the answer right away. it's due in a few hours and i have other stuff to do

OpenStudy (anonymous):

is it C? @Directrix

OpenStudy (anonymous):

@Callisto

OpenStudy (triciaal):

@iambatman stop for just one sec people don't want to be rushed for your agenda. we are here for people who want to learn. if you at least try then you won't be so overwhelmed. We are not here to give answers. we are here to help you to get to the answer. Don't be so ungrateful!

OpenStudy (anonymous):

WHO SAID I WAS UNGRATEFUL?!!? I DON'T GET THIS STUFF AND I JUST ASKED IF C WAS THE ANSWER!!

OpenStudy (anonymous):

did you get the answer? @iamabarbiegirl